# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v1

@@ 1. Introduction
The area of formal verification aims to prove the correctness of a system with respect to a specification. While the formal verification process can provide guarantees on the behavior of a system, such a result alone does not tell much about the inner workings of the system. To give some additional insight, counterexamples, invariants or related certificates as a form of verifiable justification that a system does or does not behave according to a specification have been studied extensively (see e.g., [MP95, CGP99, Nam01]). These kinds of certificates, however, do not allow us to understand the system behavior to a full extend. This establishes a desideratum for a more comprehensive understanding of why a system satisfies or violates a specification. Explications of the system are needed to assess how different components influence its behavior and performance. It is widely believed that the analysis of causal relations between occurring events during the execution of a system constitute a strong tool to form such an understanding.
@@
A lot of systems are to a certain extend influenced by probabilistic events. Thus, a branch of formal methods is studying probabilistic models such as Markov chains (MCs) which are purely probabilistic or Markov decision processes (MDPs) which combine non-determinism and probabilism. This gives rise to another approach to the concept of causality in a probabilistic setting, since the statement of counterfactual reasoning can be interpreted more gently in a probabilistic setting: Instead of saying “the effect would not have occurred, if the cause had not happened”, we can say “the probability of the effect would have been lower if the cause would not have occurred”. This interpretation leads to the widely accepted probability-raising principle which also has its roots in philosophy [Rei56, Sup70, Eel91, Hit16] and has been refined by Pearl [Pea09] for causal and probabilistic reasoning in intelligent systems. The different notions of probability-raising cause-effect relations discussed in the literature share the following two main principles:
@@
In this work we build mathematical and algorithmic foundations for probabilistic causation in Markov decision processes based on the principles (C1) and (C2). In the setting where the effect is represented as a set of terminal states, we introduce strict and global probability-raising cause sets in MDPs (Section 3). Algorithms are provided to check whether given cause and effect sets satisfy (one of) the probability-raising conditions (Section 4.1 and 4.2) and to check the existence of a cause set for a given effect set (Section 4.1). In order to evaluate the coverage properties of a cause, we subsequently introduce the above-mentioned quality measures (Section 5.1). We give algorithms for computing these values for given cause-effect sets (Section 5.2) and characterize the computational complexity of finding optimal cause sets with respect to the different measures (Section 5.3). We then extend the setting to $\omega$-regular effects (Section 6.1), and evaluate how established properties transfer to this setting. We observe that in this extension SPR causes can be viewed as a collection of GPR causes. Finally we discuss the case where causes are also path properties, namely, $\omega$-regular co-safety properties and investigate how this more general perspective affects cause-effect relations (Section 6.2). Here, the class of potential cause candidates is greatly increased, which especially affects finding optimal causes. Table 1 summarizes our complexity results.
@@
Previous work in the direction of probabilistic causation in stochastic operational models has mainly concentrated on Markov chains. Kleinberg [KM09, Kle12] introduced prima facie causes in finite Markov chains where both causes and effects are formalized as PCTL state formulae, and thus they can be seen as sets of states as in our approach. The correspondence of Kleinberg’s PCTL constraints for prima facie causes and the strict probability-raising condition formalized using conditional probabilities has been worked out in the survey article [BDF+21]. Our notion of strict probability-raising causes interpreted in Markov chains corresponds to Kleinberg’s prima facie causes with the exception of a minimality condition forbidding redundant elements in our definition. Ábrahám et al [ÁB18] introduces a hyperlogic for Markov chains and gives a formalization of probabilistic causation in Markov chains as a hyperproperty, which is consistent with Kleinberg’s prima facie causes, and with strict probability-raising causes up to minimality. Cause-effect relations in Markov chains where effects are $\omega$-regular properties and the causes are sets of paths has been introduced in [ZPF+22]. It relies on a strict probability-raising condition, but uses a probability threshold $p$ instead of directly requiring probability-raising. Therefore, [ZPF+22] permits a non-strict inequality in the PR condition with the consequence that causes always exist, which is not the case for our notions. However, a minimal good prefix of a co-safety strict probability-raising cause in a Markov chains corresponds to a probability-raising path in [ZPF+22].
The survey article [BDF+21] introduces notions of global probability-raising causes for Markov chains, where causes and effects can be path properties. [BDF+21]’s notion of reachability causes in Markov chains directly corresponds to our notion GPR causes, the only difference being that [BDF+21] deals with a relaxed minimality condition and requires that the cause set is reachable without visiting an effect state before. The latter is inherent in our approach as we suppose that all states are reachable and the effect states are terminal. On the other hand if we restrict [BDF+21]’s notion of global PR-cause to $\omega$-regular effects and co-safety causes, this corresponds to our notion co-safety GPR cause with the exception of the minimality condition. The same can be said to the correspondence of local PR-causes from [BDF+21] and co-safety SPR causes.
@@ 2. Preliminaries
Throughout the paper, we will assume some familiarity with basic concepts of Markov decision processes. Here, we present a brief summary of the notations used in the paper. For more details, we refer to [Put94, BK08, Kal20].
@@ 2.1. Markov decision processes
A Markov decision process (MDP) is a tuple $\mathcal{M}=(S,\mathit{Act},P,\mathsf{init})$ where $S$ is a finite set of states, $\mathit{Act}$ a finite set of actions, $\mathsf{init}\in S$ the initial state and $P:S\times\mathit{Act}\times S\to[0,1]$ the transition probability function such that $\sum_{t\in S}P(s,\alpha,t)\in\{0,1\}$ for all states $s\in S$ and actions $\alpha\in\mathit{Act}$.
For $\alpha\in\mathit{Act}$ and $U\subseteq S$, $P(s,\alpha,U)$ is a shortform notation for $\sum_{u\in U}P(s,\alpha,u)$. An action $\alpha$ is enabled in state $s\in S$ if $\sum_{t\in S}P(s,\alpha,t)=1$. We define $\mathit{Act}(s)=\{\alpha\in\mathit{Act}\mid\alpha\text{ is enabled in }s\}$. A state $t$ is terminal if $\mathit{Act}(t)=\emptyset$. A Markov chain (MC) is a special case of an MDP where $\mathit{Act}$ is a singleton (we then write $P(s,u)$ rather than $P(s,\alpha,u)$). A path in an MDP $\mathcal{M}$ is a (finite or infinite) alternating sequence $\pi=s_{0}\,\alpha_{0}\,s_{1}\,\alpha_{1}\,s_{2}\dots\in(S\times\mathit{Act})^{*}\times S\cup(S\times\mathit{Act})^{\omega}$ such that $P(s_{i},\alpha_{i},s_{i+1})>0$ for all indices $i$. A path is called maximal if it is infinite or finite and ends in a terminal state. If $\pi$ is a finite path in $\mathcal{M}$ then $\mathit{last}(\pi)$ denotes the last state of $\pi$. That is, if $\pi=s_{0}\,\alpha_{0}\ldots\alpha_{n-1}\,s_{n}$ then $\mathit{last}(\pi)=s_{n}$.
@@
The scheduler $\mathfrak{S}$ of $\mathcal{M}$ induces a (possibly infinite) Markov chain. We write $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}$ for the standard probability measure on measurable sets of maximal paths in the Markov chain induced by $\mathfrak{S}$ with initial state $s$. If $\varphi$ is a measurable set of maximal paths, then $\mathrm{Pr}^{\max}_{\mathcal{M},s}(\varphi)$ and $\mathrm{Pr}^{\min}_{\mathcal{M},s}(\varphi)$ denote the supremum resp. infimum of the probabilities for $\varphi$ under all schedulers. We use the abbreviation $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},\mathsf{init}}$ and notations $\mathrm{Pr}^{\max}_{\mathcal{M}}$ and $\mathrm{Pr}^{\min}_{\mathcal{M}}$ for extremal probabilities. Analogous notations will be used for expectations. So, if $f$ is a random variable, then, e.g., $\mathrm{E}^{\mathfrak{S}}_{\mathcal{M}}(f)$ denotes the expectation of $f$ under $\mathfrak{S}$ and $\mathrm{E}^{\max}_{\mathcal{M}}(f)$ its supremum over all schedulers. We use LTL-like temporal modalities such as $\Diamond$ (eventually) and $\mathrm{U}\,$ (until) to denote path properties. For $X,T\subseteq S$ the formula $X\,\mathrm{U}\,T$ is satisfied by paths $\pi=s_{0}s_{1}\dots$ such that there exists $j\geq 0$ such that for all $i<j:s_{i}\in X$ and $s_{j}\in T$ and $\Diamond T=S\,\mathrm{U}\,T$. It is well-known that $\mathrm{Pr}^{\min}_{\mathcal{M}}(X\,\mathrm{U}\,T)$ and $\mathrm{Pr}^{\max}_{\mathcal{M}}(X\,\mathrm{U}\,T)$ and corresponding optimal MD-schedulers are computable in polynomial time.
If $s\in S$ and $\alpha\in\mathit{Act}(s)$, then $(s,\alpha)$ is said to be a state-action pair of $\mathcal{M}$. An end component (EC) of an MDP $\mathcal{M}$ is a strongly connected sub-MDP containing at least one state-action pair. ECs will be often identified with the set of their state-action pairs. An EC $\mathcal{E}$ is called maximal (abbreviated MEC) if there is no proper superset $\mathcal{E}^{\prime}$ of (the set of state-action pairs of) $\mathcal{E}$ which is an EC.
@@ 2.2. MR-scheduler in MDPs without ECs
As a consequence we also have the following.
@@ 2.3. MEC-quotient
We recall the definition of the MEC-quotient, which is a standard concept for the analysis of MDPs [dA97]. More concretely, we use a modified version with an additional trap state as in [BBD+18] that serves to mimic behaviors inside an end component of the original MDP.
@@
[MEC-quotient of an MDP] Let $\mathcal{M}=(S,\mathit{Act},P,\mathsf{init})$ be an MDP with end components. Let $\mathcal{E}_{1},\ldots,\mathcal{E}_{k}$ be the MECs of $\mathcal{M}$. We may suppose without loss of generality that the enabled actions of the states are pairwise disjoint, i.e., whenever $s_{1},s_{2}$ are states in $\mathcal{M}$ with $s_{1}\not=s_{2}$ then $\mathit{Act}_{\mathcal{M}}(s_{1})\cap\mathit{Act}_{\mathcal{M}}(s_{2})=\varnothing$. This permits to consider $\mathcal{E}_{i}$ as a subset of $\mathit{Act}$. Let $U_{i}$ denote the set of states that belong to $\mathcal{E}_{i}$ and let $U=U_{1}\cup\ldots\cup U_{k}$. The MEC-quotient of $\mathcal{M}$ is the MDP $\mathcal{N}=(S^{\prime},\mathit{Act}^{\prime},P^{\prime},\mathsf{init}^{\prime})$ and the function $\iota:S\to S^{\prime}$ are defined as follows.
@@
    • If $s$ is a state of $\mathcal{M}$ that does not belong to an MEC of $\mathcal{M}$ (i.e., $s\in S\cap S^{\prime}$) then then $\mathit{Act}_{\mathcal{N}}(s)=\mathit{Act}_{\mathcal{M}}(s)$ and $P^{\prime}(s,\alpha,s^{\prime})=P(s,\alpha,\iota^{-1}(s^{\prime}))$ for all $s^{\prime}\in S^{\prime}$ and $\alpha\in\mathit{Act}_{\mathcal{M}}(s)$.
@@
Given a scheduler $\mathfrak{T}$ for $\mathcal{N}$, we pick an MD-scheduler $\mathfrak{U}$ such that such that $\mathfrak{U}(u)\in\mathcal{E}_{i}$ for each $u\in U_{i}$. Then, the corresponding scheduler $\mathfrak{S}$ for $\mathcal{M}$ behaves as $\mathfrak{T}$ as long as $\mathfrak{T}$ does not choose the $\tau$-transition to $\bot$. As soon as $\mathfrak{T}$ schedules $\tau$ then $\mathfrak{S}$ behaves as $\mathfrak{U}$ from this moment on.

### v2

@@ 1. Introduction
The area of formal verification aims to prove the correctness of a system with respect to a specification. While the formal verification process can provide guarantees on the behavior of a system, such a result alone does not tell much about the inner workings of the system. To give some additional insight, counterexamples, invariants or related certificates as a form of verifiable justification that a system does or does not behave according to a specification have been studied extensively (see e.g., [MP95, CGP99, Nam01]). These kinds of certificates, however, do not allow us to understand the system behavior to a full extend. In epistemic terms, the outcome of model checking applied to a system and a specification provides knowledge that a system satisfies a specification (or not) in terms of an assertion (whether the system satisfies the specification) and a justification (certificate or counterexample) to increase the belief in the result. However, model checking usually does not provide understanding on why a system behaves in a certain way. This establishes a desideratum for a more comprehensive understanding of why a system satisfies or violates a specification. Explications of the system are needed to assess how different components influence its behavior and performance. Causal relations between occurring events during the execution of a system can constitute a strong tool to form such an understanding. Moreover, causality is fundamental for determining moral responsibility [CH04, BvH12] or legal accountability [FHJ+11], and ultimately fosters user acceptance through an increased level of transparency [Mil17].
@@
A lot of systems are to a certain extend influenced by probabilistic events. Thus, a branch of formal methods is studying probabilistic models such as Markov chains (MCs) which are purely probabilistic or Markov decision processes (MDPs) which combine non-determinism and probabilistic choice. This gives rise to another approach to the concept of causality in a probabilistic setting, since the statement of counterfactual reasoning can be interpreted more gently in a probabilistic setting: Instead of saying “the effect would not have occurred, if the cause had not happened”, we can say “the probability of the effect would have been lower if the cause would not have occurred”. This interpretation leads to the widely accepted probability-raising principle which also has its roots in philosophy [Rei56, Sup70, Eel91, Hit16] and has been refined by Pearl [Pea09] for causal and probabilistic reasoning in intelligent systems. The different notions of probability-raising cause-effect relations discussed in the literature share the following two main principles:
@@
In this work we build mathematical and algorithmic foundations for probabilistic causation in Markov decision processes based on the principles (C1) and (C2). In the setting where the effect is represented as a set of terminal states, we introduce strict and global probability-raising cause sets in MDPs (Section 3). Algorithms are provided to check whether given cause and effect sets satisfy (one of) the probability-raising conditions (Section 4.1 and 4.2) and to check the existence of a cause set for a given effect set (Section 4.1). In order to evaluate the coverage properties of a cause, we subsequently introduce the above-mentioned quality measures (Section 5.1). We give algorithms for computing these values for given cause-effect sets (Section 5.2) and characterize the computational complexity of finding optimal cause sets with respect to the different measures (Section 5.3). We then extend the setting to $\omega$-regular effects (Section 6.1), and evaluate how established properties transfer to this setting. We observe that in this extension SPR causes can be viewed as a collection of GPR causes. Finally we discuss the case where causes are also path properties, namely, $\omega$-regular co-safety properties and investigate how this more general perspective affects cause-effect relations (Section 6.2). Here, the class of potential cause candidates is greatly increased. Table 1 summarizes our complexity results.
@@
Previous work in the direction of probabilistic causation in stochastic operational models has mainly concentrated on Markov chains. Kleinberg [KM09, Kle12] introduced prima facie causes in finite Markov chains where both causes and effects are formalized as PCTL state formulae, and thus they can be seen as sets of states as in our approach. The correspondence of Kleinberg’s PCTL constraints for prima facie causes and the strict probability-raising condition formalized using conditional probabilities has been worked out in the survey article [BDF+21]. Our notion of strict probability-raising causes interpreted in Markov chains corresponds to Kleinberg’s prima facie causes with the exception of a minimality condition forbidding redundant elements in our definition. Ábrahám et al [ÁB18] introduces a hyperlogic for Markov chains and gives a formalization of probabilistic causation in Markov chains as a hyperproperty, which is consistent with Kleinberg’s prima facie causes, and with strict probability-raising causes up to minimality. Cause-effect relations in Markov chains where effects are $\omega$-regular properties and the causes are sets of paths have been introduced in [ZPF+22]. These relations rely on a strict probability-raising condition, but use a probability threshold $p$ instead of directly requiring probability-raising. Therefore, [ZPF+22] permits a non-strict inequality in the PR condition with the consequence that causes always exist, which is not the case for our notions. However, a minimal good prefix of a co-safety strict probability-raising cause in a Markov chains corresponds to a probability-raising path in [ZPF+22].
The survey article [BDF+21] introduces notions of global probability-raising causes for Markov chains, where causes and effects can be path properties. [BDF+21]’s notion of reachability causes in Markov chains directly corresponds to our notion GPR causes, the only difference being that [BDF+21] deals with a relaxed minimality condition and requires that the cause set is reachable without visiting an effect state before. The latter is inherent in our approach as we suppose that all states are reachable and the effect states are terminal. On the other hand if we restrict [BDF+21]’s notion of global PR-cause to $\omega$-regular effects and co-safety causes, this corresponds to our notion of a co-safety GPR cause with the exception of the minimality condition. The same can be said about the correspondence of local PR-causes from [BDF+21] and co-safety SPR causes.
@@ 2. Preliminaries
Throughout the paper, we will assume familiarity with basic concepts of Markov decision processes. Here, we present a brief summary of the notations used in the paper. For more details, we refer to [Put94, BK08, Kal20].
@@ 2.1. Markov decision processes
A Markov decision process (MDP) is a tuple $\mathcal{M}=(S,\mathit{Act},P,\mathsf{init})$ where $S$ is a finite set of states, $\mathit{Act}$ a finite set of actions, $\mathsf{init}\in S$ a unique initial state and $P:S\times\mathit{Act}\times S\to[0,1]$ the transition probability function such that $\sum_{t\in S}P(s,\alpha,t)\in\{0,1\}$ for all states $s\in S$ and actions $\alpha\in\mathit{Act}$.
For $\alpha\in\mathit{Act}$ and $T\subseteq S$, $P(s,\alpha,T)$ is a shortform notation for $\sum_{t\in t}P(s,\alpha,t)$. An action $\alpha$ is enabled in state $s\in S$ if $\sum_{t\in S}P(s,\alpha,t)=1$. We define $\mathit{Act}(s)=\{\alpha\in\mathit{Act}\mid\alpha\text{ is enabled in }s\}$. A state $t$ is terminal if $\mathit{Act}(t)=\emptyset$. A Markov chain (MC) is a special case of MDP where $\mathit{Act}$ is a singleton (we then write $P(s,t)$ rather than $P(s,\alpha,t)$). A path in an MDP $\mathcal{M}$ is a (finite or infinite) alternating sequence $\pi=s_{0}\,\alpha_{0}\,s_{1}\,\alpha_{1}\,s_{2}\dots\in(S\times\mathit{Act})^{*}\times S\cup(S\times\mathit{Act})^{\omega}$ such that $P(s_{i},\alpha_{i},s_{i+1})>0$ for all indices $i$. A path is called maximal if it is infinite or finite and ends in a terminal state. If $\pi$ is a finite path in $\mathcal{M}$ then $\mathit{last}(\pi)$ denotes the last state of $\pi$. That is, if $\pi=s_{0}\,\alpha_{0}\ldots\alpha_{n-1}\,s_{n}$ then $\mathit{last}(\pi)=s_{n}$.
@@
A scheduler $\mathfrak{S}$ of $\mathcal{M}$ induces a (possibly infinite) Markov chain. We write $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}$ for the standard probability measure on measurable sets of maximal paths in the Markov chain induced by $\mathfrak{S}$ with initial state $s$. If $\varphi$ is a measurable set of maximal paths, then $\mathrm{Pr}^{\max}_{\mathcal{M},s}(\varphi)$ and $\mathrm{Pr}^{\min}_{\mathcal{M},s}(\varphi)$ denote the supremum resp. infimum of the probabilities for $\varphi$ under all schedulers. We use the abbreviation $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},\mathsf{init}}$ and notations $\mathrm{Pr}^{\max}_{\mathcal{M}}$ and $\mathrm{Pr}^{\min}_{\mathcal{M}}$ for extremal probabilities. Analogous notations will be used for expectations. So, if $f$ is a random variable, then, e.g., $\mathrm{E}^{\mathfrak{S}}_{\mathcal{M}}(f)$ denotes the expectation of $f$ under $\mathfrak{S}$ and $\mathrm{E}^{\max}_{\mathcal{M}}(f)$ its supremum over all schedulers. We also use conditional probabilities in MDPs cf. [BKKM14, Mär20]. For two measurable sets of maximal paths $\varphi$ and $\psi$ we have
$\mathrm{Pr}^{\max}_{\mathcal{M},s}(\varphi\mid\psi)=\max_{\mathfrak{S}}\ \mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\varphi\mid\psi)=\max_{\mathfrak{S}}\ \frac{\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\varphi\wedge\psi)}{\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M},s}(\psi)}$
and $\mathrm{Pr}^{\min}_{\mathcal{M},s}(\varphi\mid\psi)$ likewise. For reachability objectives maximal (and minimal) conditional probabilities are computable in polynomial time [BKKM14].
If $s\in S$ and $\alpha\in\mathit{Act}(s)$, then $(s,\alpha)$ is said to be a state-action pair of $\mathcal{M}$. Given a scheduler $\mathfrak{S}$ for $\mathcal{M}$, the expected frequencies (i.e., expected number of occurrences in maximal paths) of state action-pairs $(s,\alpha)$, states $s\in S$ and state-sets $T\subseteq S$ under $\mathfrak{S}$ are defined by:
[…]
We use LTL-like temporal modalities such as $\Diamond$ (eventually) and $\mathrm{U}\,$ (until) to denote path properties. For $X,T\subseteq S$ the formula $X\,\mathrm{U}\,T$ is satisfied by paths $\pi=s_{0}s_{1}\dots$ such that there exists $j\geq 0$ such that for all $i<j:s_{i}\in X$ and $s_{j}\in T$ and $\Diamond T=S\,\mathrm{U}\,T$. It is well-known that $\mathrm{Pr}^{\min}_{\mathcal{M}}(X\,\mathrm{U}\,T)$ and $\mathrm{Pr}^{\max}_{\mathcal{M}}(X\,\mathrm{U}\,T)$ and corresponding optimal MD-schedulers are computable in polynomial time.
An end component (EC) of an MDP $\mathcal{M}$ is a strongly connected sub-MDP containing at least one state-action pair. ECs will be often identified with the set of their state-action pairs. An EC $\mathcal{E}$ is called maximal (abbreviated MEC) if there is no proper superset $\mathcal{E}^{\prime}$ of (the set of state-action pairs of) $\mathcal{E}$ which is an EC.
@@ 2.2. MR-scheduler in MDPs without ECs
As a consequence we can build linear combinations of scheduler in such MDPs.
@@ 2.3. MEC-quotient
We recall the definition of the MEC-quotient, which is a standard concept for the analysis of MDPs [dA97]. Intuitively, the MEC-quotient of an MDP collapses all maximal end components, ignoring the actions in the end component while keeping outgoing transitions. More concretely, we use a modified version with an additional trap state as in [BBD+18] that serves to mimic behaviors inside an end component of the original MDP.
@@
[MEC-quotient of an MDP] Let $\mathcal{M}=(S,\mathit{Act},P,\mathsf{init})$ be an MDP with end components. Let $\mathcal{E}_{1},\ldots,\mathcal{E}_{k}$ be the maximal end components (MECs) of $\mathcal{M}$. We may suppose without loss of generality that enabled actions of states are pairwise disjoint, i.e., whenever $s_{1},s_{2}$ are states in $\mathcal{M}$ with $s_{1}\not=s_{2}$ then $\mathit{Act}_{\mathcal{M}}(s_{1})\cap\mathit{Act}_{\mathcal{M}}(s_{2})=\varnothing$. This permits to consider $\mathcal{E}_{i}$ as a subset of $\mathit{Act}$. Let $U_{i}$ denote the set of states that belong to $\mathcal{E}_{i}$ and let $U=U_{1}\cup\ldots\cup U_{k}$. The MEC-quotient of $\mathcal{M}$ is the MDP $\mathcal{N}=(S^{\prime},\mathit{Act}^{\prime},P^{\prime},\mathsf{init}^{\prime})$ and the function $\iota:S\to S^{\prime}$ are defined as follows.
@@
    • If $s$ is a state of $\mathcal{M}$ that does not belong to an MEC of $\mathcal{M}$ (i.e., $s\in S\cap S^{\prime}$) then $\mathit{Act}_{\mathcal{N}}(s)=\mathit{Act}_{\mathcal{M}}(s)$ and $P^{\prime}(s,\alpha,s^{\prime})=P(s,\alpha,\iota^{-1}(s^{\prime}))$ for all $s^{\prime}\in S^{\prime}$ and $\alpha\in\mathit{Act}_{\mathcal{M}}(s)$.
@@
Given a scheduler $\mathfrak{T}$ for $\mathcal{N}$, we pick an MD-scheduler $\mathfrak{U}$ such that $\mathfrak{U}(u)\in\mathcal{E}_{i}$ for each $u\in U_{i}$. Then, the corresponding scheduler $\mathfrak{S}$ for $\mathcal{M}$ behaves as $\mathfrak{T}$ as long as $\mathfrak{T}$ does not choose the $\tau$-transition to $\bot$. As soon as $\mathfrak{T}$ schedules $\tau$ then $\mathfrak{S}$ behaves as $\mathfrak{U}$ from this moment on.